OpenStudy (anonymous):

the equation of the line that is parallel to the line y=1/2x+7 and passes through the point (-4,2). in slope/intercept form.

OpenStudy (ashwinram):

Call it y=mx+c. If it's parallel to y=1/2x+7, its slope is the same. So, slope m = 1/2. Now figure out c (intercept).